Grade 8General Physics

What is the focal length of a convex lens of focal length 30cm in contact with a concave lens of focal length 20cm? Is the system a converging or a diverging lens? Ignore thickness of the lenses?

prasanjeet kumar
12 Years agoGrade 8

1 Answer

Apoorva Arora
12 Years ago
\frac{1}{f_{eq}}=\frac{1}{f_{1}}+\frac{1}{f_{2}}
f_{1}=+30 cm
f_{2}=-20 cm
so,f_{eq}=-60 cm
the focal length is negative so the resultant lens is a diverging lens.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty